The average speed of a car is 35 h-1. How far can it travel in 45 mins


1
Expert's answer
2021-02-17T11:01:16-0500
"s=vt=35\ \dfrac{km}{h}\cdot45\ min\cdot\dfrac{1\ h}{60\ min}=26.25\ km."